# fastkqr: A Fast Algorithm for Kernel Quantile Regression

### v1

@@ Appendix C Additional Simulations
Table 3: Objective values and computation time of four kernel quantile regression solvers for simulation data with $p=100$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
Following Yuan (2006), two-dimensional data are generated based on
where $X_{1}$ and $X_{2}$ are drawn from Uniform $(0,1)$, and the error term $\epsilon$ is from the standard normal distribution.
Tables 4 presents the average objective values and computation time for the four solvers applied to Example 1, with all the runtime including both model training and tuning parameter selection through five-fold cross-validation. Our algorithm, fastkqr, consistently outperforms the other solvers in speed across all the examples. For instance, when $n=1000$, our fastkqr algorithm was more than an order of magnitude faster than kernlab, and notably, more than 400 times faster than nlm and optim. As the sample size $n$ grows, the computational advantages of our algorithm over the others become more significant. Importantly, despite the significant gains in efficiency, our algorithm continues to deliver objective values that are nearly identical to those of kernlab.
$\tau$ $n$ fastkqr kernlab nlm optim
0.1 200 obj $0.442_{(0.030)}$ $0.442_{(0.030)}$ $0.442_{(0.030)}$ $0.446_{(0.030)}$
time 3.84 11.19 331.47 676.90
500 obj $0.426_{(0.020)}$ $0.426_{(0.020)}$ $0.426_{(0.020)}$ $0.428_{(0.019)}$
time 20.71 102.93 4690.49 9756.53
1000 obj $0.421_{(0.016)}$ $0.421_{(0.016)}$ $0.421_{(0.016)}$ $0.422_{(0.015)}$
time 73.83 814.82 39483.72 $>24$h
0.5 200 obj $1.031_{(0.053)}$ $1.031_{(0.053)}$ $1.032_{(0.053)}$ $1.035_{(0.052)}$
time 3.96 11.40 322.66 701.60
500 obj $0.985_{(0.035)}$ $0.984_{(0.035)}$ $0.985_{(0.035)}$ $0.986_{(0.034)}$
time 23.92 112.55 4814.45 10202.31
1000 obj $0.960_{(0.026)}$ $0.960_{(0.026)}$ $0.960_{(0.026)}$ $0.961_{(0.026)}$
time 76.64 715.08 39854.58 $>24$h
0.9 200 obj $0.468_{(0.042)}$ $0.468_{(0.042)}$ $0.469_{(0.041)}$ $0.475_{(0.039)}$
time 3.80 9.79 298.61 704.44
500 obj $0.457_{(0.020)}$ $0.457_{(0.020)}$ $0.457_{(0.020)}$ $0.459_{(0.020)}$
time 22.93 110.13 4116.96 9816.24
1000 obj $0.445_{(0.017)}$ $0.445_{(0.017)}$ $0.446_{(0.017)}$ $0.447_{(0.017)}$
time 76.01 778.49 38489.13 $>24$h
@@
Table 4: Objective values and computation time of four kernel quantile regression solvers using Yuan (2006) with $p=2$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs, and the standard errors are presented in the parentheses.
Appendix D Benchmark Data Applications
@@
We used four benchmark data from the R packages MASS and mlbench to further compare our fastkqr with the three competitors: kernlab, nlm, and optim. The first data set crabs includes five morphometric measurements from 50 crabs of the Leptograpsus variegatus species, collected in Fremantle, Western Australia. For our analysis, categorical variables were converted to dummy variables, with carapace width as the response variable, and variable “index” was removed. The second data set GAGurine (GAG) was introduced in Section 1. The third data set, BostonHousing (BH), provides housing information in the Boston area with the median value of owner-occupied homes as the predictor variable. The last data set, geyser, records details of the “Old Faithful" geyser in the Yellowstone National Park, using eruption time as the predictor variable.
@@
For each benchmark data, we employed the five-fold cross-validation to select the optimal tuning parameter, which was then used to calculate the objective values for accuracy assessment. We also recorded the total computation time, including both the model training and parameter tuning via the five-fold cross-validation. Notably, fastkqr consistently demonstrated superior efficiency compared to kernlab. Both fastkqr and kernlab achieved the highest accuracy. The computational advantages of fastkqr become increasingly evident with larger sample sizes. For instance, when fitting the mcycle data with 133 observations, fastkqr required only half the time needed by kernlab. For the larger BostonHousing data consisting of 506 observations, fastkqr was ten times faster than cvxr.
@@
$\text{data}_{(n,p)}$ $\tau$ fastkqr kernlab nlm optim
$\text{crabs}_{(200,8)}$ 0.1 obj $0.486_{(0.069)}$ $0.486_{(0.069)}$ $0.487_{(0.069)}$ $1.800_{(0.420)}$
time 3.12 9.38 263.23 731.41
0.5 obj $1.595_{(0.176)}$ $1.595_{(0.176)}$ $1.596_{(0.175)}$ $3.210_{(1.130)}$
time 2.83 10.10 241.36 801.27
0.9 obj $0.476_{(0.051)}$ $0.476_{(0.051)}$ $0.477_{(0.051)}$ $0.994_{(0.899)}$
time 3.10 9.02 261.37 803.47
$\text{GAG}_{(314,1)}$ 0.1 obj $0.685_{(0.026)}$ $0.685_{(0.026)}$ $0.685_{(0.026)}$ $0.737_{(0.044)}$
time 11.23 55.41 1326.94 3190.88
0.5 obj $2.371_{(0.140)}$ $2.371_{(0.140)}$ $2.371_{(0.140)}$ $2.372_{(0.140)}$
time 12.62 46.02 1011.65 2234.01
0.9 obj $1.652_{(0.113)}$ $1.652_{(0.113)}$ $1.652_{(0.113)}$ $1.654_{(0.112)}$
time 11.51 49.63 1080.17 3134.22
$\text{mcycle}_{(133,1)}$ 0.1 obj $4.843_{(0.212)}$ $4.843_{(0.212)}$ $4.851_{(0.212)}$ $5.580_{(0.307)}$
time 3.90 7.55 97.60 236.85
0.5 obj $9.969_{(0.489)}$ $9.969_{(0.489)}$ $9.989_{(0.487)}$ $10.340_{(0.450)}$
time 3.28 4.45 197.67 420.41
0.9 obj $4.210_{(0.527)}$ $4.210_{(0.527)}$ $4.215_{(0.524)}$ $4.469_{(0.444)}$
time 2.66 6.75 125.48 298.21
$\text{BH}_{(506,14)}$ 0.1 obj $0.826_{(0.055)}$ $0.826_{(0.055)}$ $0.827_{(0.055)}$ $1.020_{(0.221)}$
time 16.99 161.32 7473.12 12340.06
0.5 obj $1.941_{(0.056)}$ $1.941_{(0.056)}$ $1.946_{(0.055)}$ $2.079_{(0.118)}$
time 14.13 116.88 5396.74 11370.48
0.9 obj $1.051_{(0.028)}$ $1.051_{(0.028)}$ $1.072_{(0.025)}$ $1.553_{(0.455)}$
time 16.05 152.68 5317.58 10527.89
@@
Table 5: Objective values and computation time of four KQR solvers for benchmark data with $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
@@
For the NCKQR problem, we employed the same four benchmark data to compare fastkqr algorithm with its three competitors: cvxr, nlm, and optim. For each benchmark data, we presented the average objective values and the total computation time across a range of $\lambda_{2}$ values and five quantile levels: 0.1, 0.3, 0.5, 0.7, and 0.9. Table 6 show that our algorithm consistently delivers the best accuracy and the computation is the fastest. In all the examples, fastkqr consistently outperforms the other competitors, being at least 80 times faster.
@@
$\text{crabs}_{(200,8)}$ obj $0.921_{(0.023)}$ $1.051_{(0.593)}$ $7.070_{(10.791)}$ $12.280_{(6.382)}$
time 77.98 6125.56 6590.67 13067.35
$\text{GAG}_{(314,1)}$ obj $5.067_{(0.083)}$ $5.407_{(1.052)}$ $6.312_{(2.497)}$ $6.036_{(1.012)}$
time 31.03 9316.60 18614.09 39213.16
$\text{mcycle}_{(133,1)}$ obj $30.575_{(0.397)}$ $32.025_{(5.123)}$ $35.535_{(9.088)}$ $38.662_{(5.332)}$
time 10.40 6736.39 2683.09 5990.27
$\text{BH}_{(506,14)}$ obj $4.006_{(0.044)}$ $5.423_{(4.243)}$
time 321.08 17850.87 $>24$h $>24$h
Table 6: Objective values and computation time of four NCKQR solvers for benchmark data. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses. A result shown as a star “*" means the corresponding solver cannot output a solution due to some numerical issue.
Appendix E Technical Proofs
@@ E.1 Proof of Proposition 1
Thus, for any $t\in\mathbb{R}$, $40\leq{}H_{\tau,\gamma}(t)-\rho_{\tau}(t)\leq\gamma/$ 4. It follows that
$0\leq{}G^{\gamma}(b,\boldsymbol{\alpha})-G(b,\boldsymbol{\alpha})\leq\frac{\gamma}{4},\forall{}b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n},$ (32)
@@
$\begin{split}&\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}G^{\gamma}(b,\boldsymbol{\alpha})\\ &\text{subject to}\enskip{}y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in S_{0}.\end{split}$ (33)
The Lagrangian of problem (33) is
$L(b,\boldsymbol{\alpha},\xi_{i})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,{}\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}+\sum_{i\in S_{0}}\xi_{i}(b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}-y_{i}),$ (34)
where the $\xi_{i}$’s are the Lagrangian multipliers. By the optimality of $(\hat{\boldsymbol{\alpha}}^{\gamma},\hat{b}^{\gamma})$ in problem (33), we have
$\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in S_{0}}\xi_{i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in S_{0}}\xi_{i}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\,i\in S_{0},\end{array}\right.$ (35)
@@
According to the definition of $H_{\gamma,{}\tau}(t)$, one has $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\notin S_{0}$ and $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\tau-1/2{}\in{}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\in S_{0}$. Therefore, it follows from expression (35) that
$\left\{\begin{array}[]{l}\mathbf{0}\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in S_{0}}\xi_{i}\mathbf{K}_{i},\\ 0\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in S_{0}}\xi_{i}.\end{array}\right.$ (36)
@@
$\begin{split}&\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ &\text{subject to}\enskip{}y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in S_{0}.\end{split}$ (37)
Therefore, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ is the optimal solution of problem (37). This implies that
@@
The first inequality is justified because $(\hat{b},\hat{\boldsymbol{\alpha}})$ is a feasible point of problem (37). The second inequality follows since $(\hat{b},\hat{\boldsymbol{\alpha}})$ is the unconstrained minimizer of $G(b,\boldsymbol{\alpha})$. Thus, we have $G(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=G(\hat{b},\hat{\boldsymbol{\alpha}})$. Given the uniqueness of $(\hat{b},\hat{\boldsymbol{\alpha}})$, it follows that $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=(\hat{b},\hat{\boldsymbol{\alpha}})$.
E.2 Proof of Theorem 2
@@
By the optimality of $(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})$, we have $G^{\gamma}(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})\leq{}G^{\gamma}(\hat{b},\hat{\boldsymbol{\alpha}})$. It then follows from inequality (32) that
@@ E.3 Proof of Theorem 3
$\begin{split}\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}&G^{\gamma}(b,\boldsymbol{\alpha})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ \text{subject to}\enskip{}&y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in\hat{S}.\end{split}$ (38)

### v2

@@ 7 Disclosure Statement
Table S1: Objective values and computation time of seven kernel quantile regression solvers for simulation data (Friedman et al., 2010) with $p=100$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
$\tau$ $n$ fastkqr kernlab clarabel ADMM osqp nlm optim
0.1 200 obj $0.634_{(0.039)}$ $0.634_{(0.039)}$ $0.634_{(0.039)}$ $0.634_{(0.039)}$ $0.634_{(0.039)}$ $0.634_{(0.039)}$ $0.693_{(0.166)}$
time 1.41 42.73 85.60 138.07 66.34 227.24 523.11
500 obj $0.632_{(0.027)}$ $0.631_{(0.027)}$ $0.631_{(0.027)}$ $0.672_{(0.194)}$ $0.631_{(0.027)}$ $0.632_{(0.027)}$ $0.639_{(0.034)}$
time 8.35 228.24 636.40 2298.51 497.47 2987.70 7117.52
1000 obj $0.641_{(0.021)}$ $0.641_{(0.021)}$ $0.641_{(0.021)}$ $0.641_{(0.021)}$ $0.641_{(0.021)}$ $0.641_{(0.021)}$ $0.682_{(0.119)}$
time 62.32 1013.36 6601.43 11948.91 3372.22 24770.94 59033.30
0.5 200 obj $1.061_{(0.439)}$ $1.061_{(0.439)}$ $1.061_{(0.439)}$ $1.061_{(0.439)}$ $1.061_{(0.439)}$ $1.077_{(0.416)}$ $1.227_{(0.245)}$
time 1.55 59.04 79.22 412.30 73.56 261.21 525.72
500 obj $0.978_{(0.335)}$ $0.978_{(0.335)}$ $0.978_{(0.335)}$ $0.978_{(0.335)}$ $0.978_{(0.335)}$ $1.021_{(0.295)}$ $1.252_{(0.164)}$
time 8.25 283.25 563.51 2705.82 517.18 2897.41 6925.84
1000 obj $1.059_{(0.214)}$ $1.059_{(0.213)}$ $1.059_{(0.213)}$ $1.059_{(0.213)}$ $1.059_{(0.213)}$ $1.103_{(0.184)}$ $1.305_{(0.134)}$
time 62.89 1181.56 4968.14 13406.95 $3$.52 27473.93 59327.84
0.9 200 obj $0.610_{(0.037)}$ $0.610_{(0.037)}$ $0.610_{(0.037)}$ $0.610_{(0.037)}$ $0.610_{(0.037)}$ $0.610_{(0.037)}$ $0.617_{(0.047)}$
time 1.49 45.33 89.07 229.38 68.94 231.50 520.99
500 obj $0.638_{(0.032)}$ $0.638_{(0.032)}$ $0.638_{(0.032)}$ $0.679_{(0.183)}$ $0.638_{(0.032)}$ $0.638_{(0.032)}$ $0.645_{(0.036)}$
time 8.31 224.15 647.39 1720.98 485.26 2921.25 7018.28
1000 obj $0.639_{(0.020)}$ $0.639_{(0.019)}$ $0.639_{(0.019)}$ $0.682_{(0.201)}$ $0.639_{(0.019)}$ $0.639_{(0.019)}$ $0.668_{(0.050)}$
time 61.56 1017.84 5840.61 10796.91 3282.25 25235.31 58794.06
@@
Table S2: Objective values and computation time of seven kernel quantile regression solvers for simulation data (Friedman et al., 2010) with $p=5000$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
Appendix S6 Benchmark Data Applications
@@
We used four benchmark data from the R packages MASS and mlbench to further compare our fastkqr with the six competitors: kernlab, clarabel, ADMM, osqp, nlm, and optim. The first data set crabs includes five morphometric measurements from 50 crabs of the Leptograpsus variegatus species, collected in Fremantle, Western Australia. For our analysis, categorical variables were converted to dummy variables, with carapace width as the response variable, and variable “index” was removed. The second data set GAGurine (GAG) was introduced in Section 1. The third data set, BostonHousing (BH), provides housing information in the Boston area with the median value of owner-occupied homes as the predictor variable. The last data set, geyser, records details of the “Old Faithful" geyser in the Yellowstone National Park, using eruption time as the predictor variable.
@@
For each benchmark data, we employed the five-fold cross-validation to select the optimal tuning parameter, which was then used to calculate the objective values for accuracy assessment. We also recorded the total computation time, including both the model training and parameter tuning via the five-fold cross-validation. All computations were carried out on an Apple M1 (16GB) processor. Notably, fastkqr consistently demonstrated superior efficiency compared to the other six solvers. In addition, fastkqr, kernlab, clarabel and osqp achieve the highest accuracy. The computational advantages of fastkqr become increasingly evident with larger sample sizes. For instance, when fitting the mcycle data with 133 observations, fastkqr required only half the time needed by kernlab. For the larger BostonHousing data consisting of 506 observations, fastkqr was ten times faster than kernlab.
@@
$\text{data}_{(n,p)}$ $\tau$ fastkqr kernlab clarabel ADMM osqp nlm optim
$\text{crabs}_{(200,8)}$ 0.1 obj $0.204_{(0.034)}$ $0.204_{(0.034)}$ $0.204_{(0.034)}$ $0.204_{(0.034)}$ $0.205_{(0.034)}$ $0.208_{(0.034)}$ $1.091_{(0.292)}$
time 1.54 4.39 23.68 116.94 7.67 224.93 458.51
0.5 obj $0.267_{(0.049)}$ $0.267_{(0.049)}$ $0.267_{(0.049)}$ $0.267_{(0.049)}$ $0.268_{(0.049)}$ $0.285_{(0.047)}$ $1.668_{(0.828)}$
time 1.45 4.83 19.15 214.45 6.95 218.69 453.75
0.9 obj $0.235_{(0.038)}$ $0.235_{(0.038)}$ $0.235_{(0.038)}$ $0.235_{(0.038)}$ $0.236_{(0.038)}$ $0.238_{(0.038)}$ $0.949_{(0.495)}$
time 1.45 5.22 23.20 140.45 8.50 220.64 455.58
$\text{GAG}_{(314,1)}$ 0.1 obj $0.540_{(0.013)}$ $0.540_{(0.013)}$ $0.540_{(0.013)}$ $0.540_{(0.013)}$ $0.540_{(0.013)}$ $0.540_{(0.013)}$ $0.578_{(0.029)}$
time 4.77 24.51 65.32 845.47 32.98 639.42 1634.79
0.5 obj $1.460_{(0.047)}$ $1.460_{(0.047)}$ $1.460_{(0.047)}$ $1.460_{(0.047)}$ $1.460_{(0.047)}$ $1.461_{(0.047)}$ $1.500_{(0.039)}$
time 5.21 19.90 52.99 895.03 28.07 720.77 1631.48
0.9 obj $0.235_{(0.038)}$ $0.235_{(0.038)}$ $0.235_{(0.038)}$ $0.235_{(0.038)}$ $0.236_{(0.038)}$ $0.238_{(0.038)}$ $0.949_{(0.495)}$
time 4.72 21.69 72.08 729.15 33.98 562.52 1655.45
$\text{mcycle}_{(133,1)}$ 0.1 obj $4.177_{(0.259)}$ $4.177_{(0.259)}$ $4.177_{(0.259)}$ $4.177_{(0.259)}$ $4.180_{(0.259)}$ $4.196_{(0.254)}$ $5.056_{(0.384)}$
time 1.74 4.37 11.29 92.50 3.92 65.40 163.23
0.5 obj $8.565_{(0.534)}$ $8.565_{(0.534)}$ $8.565_{(0.534)}$ $8.565_{(0.534)}$ $8.567_{(0.533)}$ $8.620_{(0.527)}$ $9.254_{(0.557)}$
time 1.56 2.45 8.63 68.33 2.71 77.05 163.75
0.9 obj $3.773_{(0.141)}$ $3.773_{(0.141)}$ $3.773_{(0.141)}$ $3.773_{(0.141)}$ $3.775_{(0.141)}$ $3.781_{(0.140)}$ $4.137_{(0.250)}$
time 1.32 4.65 10.44 91.25 3.93 66.97 162.88
$\text{BH}_{(506,14)}$ 0.1 obj $0.490_{(0.020)}$ $0.490_{(0.020)}$ $0.490_{(0.020)}$ $0.490_{(0.020)}$ $0.490_{(0.019)}$ $0.498_{(0.019)}$ $0.700_{(0.056)}$
time 6.17 70.67 360.05 2427.07 136.88 3425.43 6988.44
0.5 obj $1.022_{(0.013)}$ $1.022_{(0.013)}$ $1.022_{(0.013)}$ $1.022_{(0.013)}$ $1.023_{(0.013)}$ $1.165_{(0.058)}$ $1.740_{(0.547)}$
time 5.92 58.45 302.34 2657.78 90.87 3503.75 6963.22
0.9 obj $0.671_{(0.002)}$ $0.671_{(0.002)}$ $0.671_{(0.002)}$ $0.671_{(0.002)}$ $0.671_{(0.002)}$ $0.761_{(0.010)}$ $1.438_{(0.377)}$
time 6.28 79.92 405.05 1150.37 135.20 3434.75 7038.91
@@
Table S3: Objective values of seven KQR solvers for benchmark data with $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
@@
For the NCKQR problem, we employed the same four benchmark data to compare fastkqr algorithm with its three competitors: cvxr, nlm, and optim. For each benchmark data, we presented the average objective values and the total computation time across a range of $\lambda_{2}$ values and five quantile levels: 0.1, 0.3, 0.5, 0.7, and 0.9. Table S4 show that our algorithm consistently delivers the best accuracy and the computation is the fastest. In all the examples, fastkqr consistently outperforms the other competitors, being at least 80 times faster.
@@
$\text{crabs}_{(200,8)}$ obj $0.920_{(0.005)}$ $1.053_{(0.024)}$ $7.071_{(0.317)}$ $12.264_{(0.441)}$
time 13.41 1137.84 2936.02 6420.14
$\text{GAG}_{(314,1)}$ obj $5.082_{(0.083)}$ $5.400_{(0.083)}$ $6.345_{(0.180)}$ $6.051_{(0.181)}$
time 5.82 1433.54 10385.86 22994.12
$\text{mcycle}_{(133,1)}$ obj $30.607_{(0.476)}$ $31.849_{(0.973)}$ $35.614_{(0.936)}$ $38.751_{(2.107)}$
time 2.29 1063.62 1043.04 2269.04
$\text{BH}_{(506,14)}$ obj $4.001_{(0.032)}$ $5.414_{(0.168)}$ $11.025_{(0.388)}$
time 28.87 2856.31 39887.25 >24>24> 24h
Table S4: Objective values and computation time of four NCKQR solvers for benchmark data. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses. A result shown as a star “*" means the corresponding solver cannot output a solution due to some numerical issue.
Appendix S7 Technical Proofs
@@
Thus, for any $t\in\mathbb{R}$, $0\leq{}H_{\tau,\gamma}(t)-\rho_{\tau}(t)\leq\gamma/4$. It follows that
$0\leq{}G^{\gamma}(b,\boldsymbol{\alpha})-G(b,\boldsymbol{\alpha})\leq\frac{\gamma}{4},\forall{}b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n},$ (S34)
@@
$\begin{split}&\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}G^{\gamma}(b,\boldsymbol{\alpha})\\ &\text{subject to}\enskip{}y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in S_{0}.\end{split}$ (S35)
The Lagrangian of problem (S35) is
$L(b,\boldsymbol{\alpha},\xi_{i})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,{}\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}+\sum_{i\in S_{0}}\xi_{i}(b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}-y_{i}),$ (S36)
where the $\xi_{i}$’s are the Lagrangian multipliers. By the optimality of $(\hat{\boldsymbol{\alpha}}^{\gamma},\hat{b}^{\gamma})$ in problem (S35), we have
$\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in S_{0}}\xi_{i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in S_{0}}\xi_{i}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\,i\in S_{0},\end{array}\right.$ (S37)
@@
According to the definition of $H_{\gamma,{}\tau}(t)$, one has $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\notin S_{0}$ and $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\tau-1/2{}\in{}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\in S_{0}$. Therefore, it follows from expression (S37) that
$\left\{\begin{array}[]{l}\mathbf{0}\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in S_{0}}\xi_{i}\mathbf{K}_{i},\\ 0\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in S_{0}}\xi_{i}.\end{array}\right.$ (S38)
@@
$\begin{split}&\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ &\text{subject to}\enskip{}y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in S_{0}.\end{split}$ (S39)
Therefore, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ is the optimal solution of problem (S39). This implies that
@@
The first inequality is justified because $(\hat{b},\hat{\boldsymbol{\alpha}})$ is a feasible point of problem (S39). The second inequality follows since $(\hat{b},\hat{\boldsymbol{\alpha}})$ is the unconstrained minimizer of $G(b,\boldsymbol{\alpha})$. Thus, we have $G(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=G(\hat{b},\hat{\boldsymbol{\alpha}})$. Given the uniqueness of $(\hat{b},\hat{\boldsymbol{\alpha}})$, it follows that $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=(\hat{b},\hat{\boldsymbol{\alpha}})$.
S7.2 Proof of Theorem 2
@@
By the optimality of $(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})$, we have $G^{\gamma}(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})\leq{}G^{\gamma}(\hat{b},\hat{\boldsymbol{\alpha}})$. It then follows from inequality (S34) that
@@
$\begin{split}\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}&G^{\gamma}(b,\boldsymbol{\alpha})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ \text{subject to}\enskip{}&y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in\hat{S}.\end{split}$ (S40)